Determine whether each of the following statements about Fibonacci numbers is true or false.

a.) 𝐹𝑛+1 = 𝐹𝑛2 + 𝐹(𝑛−1)3 for every n ≥ 2

b.) 𝐹𝑛 = 2𝐹𝑛+1 for every n ≥ 2



1
Expert's answer
2021-03-12T06:54:27-0500

Fibonacci series : 0,1,1,2,3,5,8,13...respective to f1,f2,f3,f4,f5


Fn+1=Fn2 + F(n-1)3              n≥2

Simplify the equation

Fn+1= Fn2 +F3n-3

Substitute n with 2

F(2+1)=F(2*2)+F([3*2]-3)

We get

F3=F4+F3 ...

which is false

 

b)

Fn=2Fn+1    where n≥2

Fn =2Fn+1

F2=2F2+1

F2=2F3

Which is false